Dearest tutors.,,, help me with my problem pls!!!
3. One number is 5 greater than another number. If the sum of their squares is 5 times the square of the smaller number, what are the numbers?
Plssss have mercy!!!

Answer by stanbon(75887)

One number is 5 greater than another number. If the sum of their squares is 5 times the square of the smaller number, what are the numbers?
---------
Let one of the numbers be "x"
The other is "x+5".
EQUATION:
x^2 + (x+5)^2 = 5x^2
x^2+x^2+10x+25 = 5x^2
-3x^2+10x+25=0
Use the quadratic formula to get:
x=[-10+-sqrt(100+4(3)(25)]/-6
x=[-10-20]/-6 or x=[-10+20]/-6
x=5 or x=-5/3
Cheers,
Stan H.